# Artificial Intelligence Expands Scientists’ Impact but Contracts Science’s Focus

### v2

@@ Abstract
Recent decades have witnessed unprecedented development in Artificial Intelligence (AI) to accelerate scientific discovery [1]. Alongside two recent AI-oriented Nobel prizes [2, 3, 4, 5, 6, 7, 8, 9], these trends establish the role of AI in science [10]. This advancement raises questions about the potential influences of AI on scientists and science as a whole, and highlights a potential conflict between individual and collective benefits [11]. To evaluate these concerns, we used a highly accurate pretrained language model to identify AI-augmented research, with an F1-score of 0.875 in validation against expert-labeled data. Using a dataset of 41.3 million research papers across the natural science and covering distinct eras of AI, here we show an accelerated adoption of AI among scientists and consistent professional advantages associated with AI use, but a collective narrowing of scientific concerns and a decrease in follow-on scientist engagement. Scientists who engage in AI-augmented research publish 3.02 times more papers, receive 4.84 times more citations, and become research project leaders 1.37 years (15.75%) earlier than those who do not. By contrast, AI research shrinks the collective volume of scientific topics studied by 4.63% and decreases scientist’s engagement with one another by 22.00% when they build upon AI-augmented work. In this way, AI adoption in science presents a seeming paradox—an expansion of individual scientists’ impact but a contraction in collective science’s reach—as AI-augmented work moves collectively toward areas richest in data. With reduced follow-on scientific engagement, AI appears to automate established fields rather than explore new ones, highlighting a tension between personal advancement and collective scientific progress.
@@ Results
Increasing prevalence of AI adoption in science
@@
AI enlarges the impact of individual papers and enhances the career of individual scientists
@@
AI adoption is associated with a contraction in knowledge focus among scientific fields
@@
Reduced follow-on engagement and more overlapping works in AI research
@@
In order to analyze mechanisms underlying the conflict between the growing influence of individual papers and researchers and the narrowing of domain knowledge within AI research, we examine the relationship between the articles that cite AI and non-AI work. We first examine the knowledge extent of “paper families”, i.e., a focal paper and its follow-on citations, which measures the size of the space covered by research derived from each original paper (Fig. 4a and Methods M9). Results show that the knowledge extent of AI papers’ citation families is on average 3.46% more expanded compared to non-AI papers’ (, and in t-test on 30 out of 32 pairs of data). Therefore, the contraction of knowledge space in AI research is not attributable to the narrowing of knowledge space that can be derived from each original work of research.
@@
To further investigate, we examine the relationships between papers by measuring the degree of follow-on paper engagement, namely how frequently citations of the same original paper cite each other (Fig. 4b and Methods M10). Results demonstrate AI research spawns 22.00% less follow-on engagement (, and in t-test on any discipline), suggesting that AI papers tend to only concentrate on the original paper, rather than forming densely interactions among each other, where the latter is the characteristic of emerging fields [45]. This results in a star-like structure around specific popular research topics, rather than a network of emergent and interconnected research works. Further evidence of this concentration is found in the Matthew effect [46] among AI paper citations across different fields (Fig. 4c and Extended Data Fig. 10). In AI research, a small number of superstar papers dominate the field, with 22.20% of top papers receiving 80% of the citations and the top 54.14% receiving 95% of citations. This unequal distribution leads to a GINI coefficient of 0.754 in citation patterns surrounding AI research, higher than 0.690 for non-AI papers (, and in t-test), signaling a disparity in recognition.
@@
To further analyze the impact of reduced follow-on engagement, we sample 590,325,130 pairs of papers, where each pair cites the same original work. Among these, 51,723,984 pairs not only cite the same original work but also cite each other (engaged), while the remaining pairs do not cite each other (disengaged). We examine distances between these pairs of papers within the 768-dimensional vector space (Fig. 4d) and find that median distance between paper pairs disengaged from one another tends to be 18.11% larger than between paper pairs engaged with each other (, and in median-test on any discipline). In contrast, the closest disengaged paper pairs are 76.51% closer to one another than the closest paper pairs engaged with one another. Taken together, a pair of disengaged papers commonly focus on less related topics and lie farther apart in the embedding space. Occasionally, however, due to the lack of reciprocal engagement, it is possible that mutually-unaware papers lie very close to each other, which indicates more overlapping research. These findings suggest that AI in science has become more concentrated around popular research topics that become “lonely crowds” with reduced interaction among papers, linking to more overlapping research and a contraction in knowledge extent and diversity across science.
@@ 4.3 Robustness of distance calculations in the high-dimensional paper embedding space
Acknowledgments
Y.L. received support from National Natural Science Foundation of China #U23B2030. J.E. received support from NSF #2404109 and AFOSR #FA9550-19-1-0354.
Author contributions
F.X., Y.L. and J.E. jointly launched this research and designed the research outline. Q.H. analyzed the data and prepared the figures. All authors jointly participated in writing and revising the manuscript.
Competing interests
J.E. has a commercial affiliation with Google, but Google had no role in the design and analysis of this study. The authors declare no other competing interests.
Additional information
Supplementary information is available for this manuscript.

### v3

# Artificial Intelligence Tools Expand Scientists’ Impact but Contract Science’s Focus

@@ Abstract
Development in Artificial Intelligence (AI) has accelerated scientific discovery [1]. Alongside recent AI-oriented Nobel prizes [2, 3, 4, 5, 6, 7, 8, 9], these trends establish the role of AI tools in science [10]. This advancement raises questions about the potential influences of AI tools on scientists and science as a whole, and highlights a potential conflict between individual and collective benefits [11]. To evaluate, we used a pretrained language model to identify AI-augmented research, with an F1-score of 0.875 in validation against expert-labeled data. Using a dataset of 41.3 million research papers across natural science and covering distinct eras of AI, here we show an accelerated adoption of AI tools among scientists and consistent professional advantages associated with AI usage, but a collective narrowing of scientific focus. Scientists who engage in AI-augmented research publish 3.02 times more papers, receive 4.84 times more citations, and become research project leaders 1.37 years earlier than those who do not. By contrast, AI adoption shrinks the collective volume of scientific topics studied by 4.63% and decreases scientist’s engagement with one another by 22.00%. Thereby, AI adoption in science presents a seeming paradox—an expansion of individual scientists’ impact but a contraction in collective science’s reach—as AI-augmented work moves collectively toward areas richest in data. With reduced follow-on engagement, AI tools appear to automate established fields rather than explore new ones, highlighting a tension between personal advancement and collective scientific progress.
@@ Results
Increasing prevalence of AI in science
@@
AI enhances individual scientists
@@
AI contracts science’s focus
@@
AI reduces scientific engagement
@@
In order to analyze mechanisms underlying the conflict between the growing influence of individual papers and researchers and the narrowing of domain knowledge within AI research, we examine the relationship between articles that cite AI and non-AI work. We first examine the knowledge extent of “paper families”, i.e., a focal paper and its follow-on citations, which measures the size of the space covered by research derived from each original paper (Fig. 4a and Methods M9). Results show that the knowledge extent of AI papers’ citation families is on average 3.46% more expanded compared to non-AI papers’ (, and in t-test on 30 out of 32 pairs of data). Therefore, the contraction of knowledge space in AI research is not attributable to the narrowing of knowledge space that can be derived from each original research work.
@@
To further investigate, we examine relationships between papers by measuring the degree of follow-on paper engagement, namely how frequently citations of the same original paper cite each other (Fig. 4b and Methods M10). Results demonstrate AI research spawns 22.00% less follow-on engagement (, and in t-test on any discipline), suggesting that AI papers tend to only concentrate on the original paper, rather than forming dense interactions among each other, which is the characteristic of emerging fields [45]. This results in a star-like structure around specific popular research topics, rather than a network of emergent and interconnected research works. Further evidence of this concentration is found in the Matthew effect [46] among AI paper citations across different fields (Fig. 4c and Extended Data Fig. 10). In AI research, a small number of superstar papers dominate the field, with 22.20% of top papers receiving 80% of the citations and the top 54.14% receiving 95% of citations. This unequal distribution leads to a GINI coefficient of 0.754 in citation patterns surrounding AI research, higher than 0.690 for non-AI papers (, and in t-test), signaling a disparity in recognition.
@@
To further analyze the impact of reduced follow-on engagement, we sample 590,325,130 pairs of papers, where each pair cites the same original work. Among these, 51,723,984 pairs not only cite the same original work but also cite each other (engaged), while the remaining pairs do not cite each other (disengaged). We examine distances between these pairs of papers within the 768-dimensional vector space (Fig. 4d) and find that median distance between paper pairs disengaged from one another tends to be 18.11% larger than between paper pairs engaged with each other. In contrast, the closest disengaged paper pairs are 76.51% closer to one another than the closest paper pairs engaged with one another. Taken together, a pair of disengaged papers commonly focus on less related topics and lie farther apart in the embedding space. Occasionally, however, due to the lack of reciprocal engagement, it is possible that mutually-unaware papers lie very close to each other, which indicates more overlapping research. These findings suggest that AI in science has become more concentrated around popular research topics that become “lonely crowds” with reduced interaction among papers, linking to more overlapping research and a contraction in knowledge extent and diversity across science.